# PRISM-Loc: a Lightweight Long-range LiDAR Localization in Urban Environments with Topological Maps

### v1

@@ Abstract
Localization in the environment is one of the crucial tasks of navigation of a mobile robot or a self-driving vehicle. For long-range routes, performing localization within a dense global lidar map in real time may be difficult, and the creation of such a map may require much memory. To this end, leveraging topological maps may be useful. In this work, we propose PRISM-Loc – a topological map-based approach for localization in large environments. The proposed approach leverages a twofold localization pipeline, which consists of global place recognition and estimation of the local pose inside the found location. For local pose estimation, we introduce an original lidar scan matching algorithm, which is based on 2D features and point-based optimization. We evaluate the proposed method on the ITLP-Campus dataset on a 3 km route, and compare it against the state-of-the-art metric map-based and place recognition-based competitors. The results of the experiments show that the proposed method outperforms its competitors both quality-wise and computationally-wise.
@@ I INTRODUCTION
Despite utilizing a graph-based structure, our approach still requires solving a rough pose estimation problem. This is achieved by matching incoming LiDAR scans against a database of precomputed topological nodes, each represented by a lightweight, image-like descriptor. This process yields a candidate location that approximates the vehicle’s global position.
In the second stage, we refine this estimation using a novel LiDAR scan matching algorithm. This algorithm, based on 2D feature extraction and point-based optimization, accurately determines the local pose within the identified candidate region.
In this study, we compare the proposed method against two state-of-the-art localization techniques using a large-scale outdoor urban dataset collected with the Clearpath Husky robot. The first method relies solely on point cloud data for localization, while the second integrates place recognition and scan matching to enhance accuracy and robustness. To evaluate the efficiency of these approaches, we analyze their computational and memory performance, comparing them with our proposed method. This assessment provides insights into the trade-offs between accuracy, efficiency, and scalability in large-scale urban environments, highlighting the advantages of our approach for real-world robotic applications.
To sum up, the contributions of this work are following:
@@ II-A Long-range Mapping
The construction of large-scale metric maps for mobile robots has been extensively studied. One of the most common mapping methods is RTAB-Map [1], which employs a unique memory management technique for large-scale mapping. However, an evaluation conducted in [2] shows that, over long distances and with a noisy positioning source, RTAB-Map’s memory consumption increases significantly (up to 8 GB for a 1 km trajectory) and produces inconsistent maps with artifacts such as ”corridor bifurcation” (see Fig. 1).
Figure 1: RTAB-Map’s map with noisy positioning on a 1 km trajectory. One corridor is mapped twice.
To mitigate memory and computational costs, metric submap-based methods such as Voxgraph [3] can be utilized. The GLIM method [4] constructs a highly accurate metric map by representing the environment as a set of submaps, which are then stitched together into a voxelized point cloud. The authors tested this method on a 2 km trajectory in a park, achieving good trajectory estimation quality in real-time. However, memory consumption was not addressed in their study.
An alternative approach for long-range, large-scale mapping is the use of topological maps instead of dense global metric maps. This strategy has been implemented in works such as [5] and [6]. In [6], a topological map is represented as a graph of rooms, where doors serve as edges, and a metric map is constructed for each room separately. In [5], a city-scale topological map is built by fusing 2D global open-source maps with 3D point clouds obtained by a robot. This method can map a 17 $km^{2}$ urban indoor-outdoor environment while consuming only 1 GB of RAM. However, neither work provides a detailed description of the localization procedure within the topological map.
The TLF method [7] constructs a topological map based on multiple runs, facilitating lifelong localization under varying daytime, weather, and dynamic conditions. Localization is performed using a stereo camera, achieving an accuracy of 0.5 m and 4 degrees over multiple runs in different environments.
In recent years, learning-based topological mapping methods have emerged. For example, [8] introduces a framework for lifelong navigation, localization in a pre-built topological map, and map updates to accommodate dynamic changes. Localization is performed using a neural network that compares the current image with reference images from the topological graph. While such methods achieve high navigation success rates using stereo or RGB-D cameras, they are typically restricted to indoor environments and short trajectories.
@@ II-B Place Recognition
Early place recognition methods relied on the bag-of-words approach, where each observation is represented as an unordered set of descriptors. In visual place recognition, one of the most common bag-of-words-based methods is DBoW2 [9]. In the context of point cloud-based place recognition, BoW3D [10] has been successfully applied for loop closure on large outdoor trajectories.
@@
For vision-based place recognition, methods such as NetVLAD [11], CosPlace [12], and MixVPR [13] use convolutional or transformer-based architectures to predict descriptors. These methods achieve a Top-1 place recognition recall above 0.85 in outdoor urban environments.
For point cloud-based place recognition, PointNetVLAD [14] extends the NetVLAD architecture to 3D data. Other methods, such as MinkLoc3D [15] and its enhancement [16], utilize voxelization and sparse convolutions. SVT-Net [17] further refines this process by incorporating attention mechanisms. These approaches achieve recall rates of 0.87–0.93 in outdoor place recognition for urban environments.
Multimodal place recognition methods have also been developed, integrating image and point cloud data to improve performance. Examples include MinkLoc++ [18] and MSSPlace [19]. In our work, we utilize MinkLoc3D for place recognition as a lightweight and real-time method that relies solely on LiDAR point clouds.
@@ II-C Pose Estimation
For vision-based pose estimation, an example is StereoScan [20], which determines the transformation using stereo image pairs.
For point cloud-based pose estimation, a common approach is Iterative Closest Point (ICP) [21], which refines the transformation between two point clouds given an initial rough estimate. The RANSAC algorithm [22] is often used to generate this initial guess. However, the RANSAC + ICP pipeline is computationally expensive for dense point clouds and performs well only when the input point clouds have significant overlap. According to [23], this pipeline achieves a recall of 0.75 for point cloud pairs with overlap above 50%, but drops to 0.53 for pairs with overlap above 25%.
Using 2D projections of point clouds (e.g., bird’s-eye views) can improve pose estimation accuracy when overlap is low while also accelerating the matching process compared to full 3D point clouds. Methods such as BVMatch [24] and its enhancement BEVPlace++ [25] introduce a density-based bird’s-eye-view extraction algorithm and a novel local descriptor, achieving a pose estimation error of 0.4 m.
In our previous work [23], we introduced a fast bird’s-eye-view matching algorithm based on ORB feature matching with iterative outlier removal. In this work, we enhance our scan matching algorithm for outdoor environments by incorporating curb detection in point clouds and point-to-distance-map optimization.
@@ III PROBLEM STATEMENT
$\widehat{p_{t}^{next}}=Loc(M;C_{t}^{next};o_{t}^{next};\widehat{p_{t-1}^{next}})$
@@
$e_{t}=||\widehat{p_{t}^{next}}-p_{t}^{next}||_{2}$ (1)
$ATE_{mean}=\frac{1}{T^{N}}\sum\limits_{t=1}^{T^{N}}e_{t}$ (2)
$ATE_{median}=Median(e_{1},\dots,e_{T^{N}})$ (3)
Also we evaluate the localization success rate as the part of time steps where the localization error is lower than certain threshold:
$SR_{loc}=\frac{1}{T^{N}}\sum\limits_{t=1}^{T^{N}}I(e_{t}<Th)$ (4)
@@ IV METHOD
IV-A Map creation
For map creation, we use the concept of a graph of locations and its implementation from PRISM-TopoMap method [23]. We build the map $M=(V,E)$, where $V=(v_{1},\dots,v_{N})$ are the locations with their ground truth poses $p_{1},\dots,p_{N}$ respectively, and $E=\{(v_{i}\in V,v_{j}\in V):\ ||p_{i}-p_{j}||<5\}$ are the edges between locations (we link each pair of locations whose observation points are located not further than 5 m from each other). As the relative poses for graph edges, we use the difference between ground truth poses of the locations. For each location, we store the 2D grid extracted from its point cloud by 2D projection, and the descriptor for place recognition. So, each location consumes only several kilobytes in disk storage, which lets us store even city-scale topological maps onboard. An example of a topological map is shown in Fig. 2.
@@ IV-B Localization
IV-B1 Overall pipeline
The proposed localization pipeline maintains current state in the graph of locations. At each step $t$, this state is represented as the location $v_{cur}^{t}$ where the robot is currently located, and the robot’s relative position $T_{cur}^{t}$ in this location. The pipeline is similar to the graph maintaining pipeline of the original PRISM-TopoMap method [23], however, it has no part of a new node addition and checks the localization results for proximity to the current state:
@@ IV-B1 Overall pipeline
  • If we moved outside $v_{cur}^{t-1}$, or our overlapping percent is low, we first try to change $v_{cur}$ to one of its neighbors in the graph (i.e. move along an edge). For this purpose, we use the approximate relative position $((T_{cur}^{t-1})^{-1}(p_{v_{cur}})^{-1}p_{v_{next}})^{-1}$ as an initial guess for scan matching. If we matched successfully to one the neighbors $v_{next}$, then we change $v_{cur}$ to it: $v_{cur}^{t}=v_{next};T_{cur}^{t}=T_{next}$. To prevent false matches, we use a ”jumping threshold” – the maximum distance between an initial guess and the found transform when the matching is considered successful.
@@
  • After all, if we could not change $v_{cur}$ by edges and by the localization results, we change its value to the nearest neighbor in the graph without transform alignment (i.e. try to move along an edge without matching): $v_{cur}^{t}=v_{next};T_{cur}^{t}=((T_{cur}^{t-1})^{-1}(p_{v_{cur}})^{-1}p_{v_{next}})^{-1}$. We do this change if $||((T_{cur}^{t-1})^{-1}(p_{v_{cur}})^{-1}p_{v_{next}})^{-1}||<5$ to stay inside the next location. Otherwise, we do not change $v_{cur}$, apply step (1), and consider that the localization is lost.
@@
$\widehat{p_{t}^{next}}=p_{v_{cur}^{t}}^{first}T_{cur}^{t}$
@@
IV-B2 Place Recognition
For place recognition, we use MinkLoc3D model [15]. For each scan from the robot, we find top-5 nearest locations in the topological map using proximity of the descriptors extracted by the MinkLoc3D model. We then identify the top-5 most similar place descriptors within the nodes of the location graph based on Euclidean distance. These five candidates are utilized in the subsequent stages of our pipeline.
IV-B3 Scan Matching
@@
To estimate the transform between the reference scan $R$ from the robot, and candidate scan $C$ from the map, we combine three techniques. First, we match the scans using the features $R_{f}$ and $C_{f}$ extracted by ORB detector [26] (for global localization) and by Harris corner feature detector [27] for matching along an edge. We find the correspondences between features using FLANN matcher [28], as in the PRISM-TopoMap method.
Also, we create distance maps ${}^{c}D,^{w}D$ from curb and wall points of the candidate scan. Within these distance maps, we optimize the distances from the curbs and walls points of the current robot’s scan $(^{c}P,^{w}P)$ to the nearest curbs and walls on the candidate scan. To do this, we perform weighted Gauss-Newton optimization of the function consisting of three components:
  1. 1.
    $PointToPointFactor(\Omega_{f},R_{f},C_{f})$) is the square distance between the feature correspondences;
  2. 2.
    $PointToDistMapFactor(\Omega_{w},^{w}P,^{w}D)$ is the square distance from each wall point of the reference scan to the nearest one in the candidate scan;
  3. 3.
    $PointToDistMapFactor(\Omega_{c},^{c}P,^{c}D)$ is the square distance from each curb point of the reference scan to the nearest one in the candidate scan;
Figure 4: Scan matching comparison: our previous algorithm (left) [23] and the proposed algorithm (right). The candidate scan is shown in gray and white, the reference scan is shown in green (wall points) and red (curb points)
To make our scan matching more robust, we optimize iteratively, removing outliers from feature correspondences. We use a ”funnel” of thresholds to filter out outliers. The procedure of the matching is described in the Algorithm 1. An example of the comparison between the proposed scan matching and the PRISM-TopoMap’s one is shown in Fig. 4.
1
Data: $R_{f}$, $C_{f}$, ${}^{w}D$, ${}^{c}D$, $\nabla^{w}D$, $\nabla^{c}D$, ${}^{w}P$, ${}^{c}P$
Result: $X\in SE(2)$ is transform $T_{C,R}$, where C and R are frames attached to candidate grid and reference grid respectively
2 $X\leftarrow\text{transform from points $R_{p}$ , $C_{p}$}$ ;
3 for $\delta\in(15,7,1,0.1)$ do
4
5      $X\leftarrow$ Optimize ($X$, [
6       PointToPointFactor($\Omega_{f}$, $R_{f}$, $C_{f}$),
7       PointToDistMapFactor($\Omega_{w}$, ${}^{w}P$, ${}^{w}D$, …),
8       PointToDistMapFactor($\Omega_{c}$, ${}^{c}P$, ${}^{c}D$, …)
9       ], $N_{iter}$ = 10) ;
10
      // Outliers filtering
11       remove pairs of point $(r,c)$ $r\in R_{p}$ $c\in C_{p}$ from $R_{p}$ $C_{p}$ if $||\tau(X,r)-c||_{2}>\delta$ ;
12
13
// optimization use only walls, if we did not match
14 if $IOU(\tau(X,R),C)<0.4$ then
15
16      $X\leftarrow$ Optimize ($X$, [
17
18      PointToDistMapFactor($\Omega_{w}$, ${}^{w}P$, ${}^{w}D$, …)
19       ], $N_{iter}$ = 20) ;
20
Algorithm 1 Scan matching optimization

### v2

@@ Abstract
We propose PRISM-Loc – a lightweight and robust approach for localization in large outdoor environments that combines a compact topological representation with a novel scan-matching and curb-detection module operating on raw LiDAR scans. The method is designed for resource-constrained platforms and emphasizes real-time performance and resilience to common urban sensing challenges. It provides accurate localization in compact topological maps using global place recognition and an original scan matching technique. Experiments on standard benchmarks and on an embedded platform demonstrate the effectiveness of our approach. Our method achieves a 99% success rate on the large-scale ITLP-Campus dataset while running at 150 ms per localization and using a 20 MB map for localization. We highlight three main contributions: (1) a compact representation for city-scale localization; (2) a novel curb detection and scan matching pipeline operating directly on raw LiDAR points; (3) a thorough evaluation of our method with performance analysis.
Figure 1: Ground truth metric map (top), metric map built by RTAB-Map (middle), and topological map aligned with a ground-truth metric map (bottom).
@@ I INTRODUCTION
Despite utilizing a graph-based structure, our approach still requires solving a rough pose estimation problem. This is achieved by matching incoming LiDAR scans against a database of precomputed topological nodes, each represented by a lightweight descriptor. This process yields a candidate location that approximates the vehicle’s global position.
In the second stage, we refine this estimation using a novel LiDAR scan matching algorithm. This algorithm, based on 2D feature extraction and point-based optimization, accurately determines the local pose within the identified candidate region. To increase accuracy of matching in urban environments, an original algorithm of curbs detection is used.
In this study, we compare the proposed method against two state-of-the-art localization techniques using three large-scale outdoor urban datasets. The first method relies solely on point cloud data for localization, while the second integrates place recognition and scan matching to enhance accuracy and robustness. To evaluate the efficiency of these approaches, we analyze their computational and memory performance, comparing them with our proposed method. This assessment provides insights into the trade-offs between accuracy, efficiency, and scalability in large-scale urban environments, highlighting the advantages of our approach for real-world robotic applications.
To sum up, the contributions of this work are as follows:
@@ II-A Long-range Mapping
The construction of large-scale metric maps for mobile robots has been extensively studied. One of the most common mapping methods is RTAB-Map [labbe2019rtab], which employs a unique memory management technique for large-scale mapping. However, an evaluation conducted in [muravyev2022evaluation] shows that, over long distances and with a noisy positioning source, RTAB-Map’s memory consumption increases significantly (up to 8 GB for a 1 km trajectory) and produces inconsistent maps with artifacts such as ”corridor bifurcation” (see Fig. 1).
To mitigate memory and computational costs, metric submap-based methods such as Voxgraph [reijgwart2019voxgraph] can be utilized. The GLIM method [koide2024glim] constructs a highly accurate metric map by representing the environment as a set of submaps, which are then stitched together into a voxelized point cloud. The authors tested this method on a 2 km trajectory in a park, achieving good trajectory estimation quality in real-time. However, memory consumption was not addressed in their study.
An alternative approach for long-range, large-scale mapping is the use of topological maps instead of dense global metric maps. This strategy has been implemented in works such as [niijima2020city] and [gomez2020hybrid]. In [gomez2020hybrid], a topological map is represented as a graph of rooms, where doors serve as edges, and a metric map is constructed for each room separately. In [niijima2020city], a city-scale topological map is built by fusing 2D global open-source maps with 3D point clouds obtained by a robot. This method can map a 17 $km^{2}$ urban indoor-outdoor environment while consuming only 1 GB of RAM. However, neither work provides a detailed description of the localization procedure within the topological map.
The TLF method [tang2019topological] constructs a topological map based on multiple runs, facilitating lifelong localization under varying daytime, weather, and dynamic conditions. Localization is performed using a stereo camera, achieving an accuracy of 0.5 m and 4 degrees over multiple runs in different environments.
In recent years, learning-based topological mapping methods have emerged. For example, [wiyatno2022lifelong] introduces a framework for lifelong navigation, localization in a pre-built topological map, and map updates to accommodate dynamic changes. Localization is performed using a neural network that compares the current image with reference images from the topological graph. While such methods achieve high navigation success rates using stereo or RGB-D cameras, they are typically restricted to indoor environments and short trajectories.
@@ II-B Place Recognition
Early place recognition methods relied on the bag-of-words approach, where each observation is represented as an unordered set of descriptors. In visual place recognition, one of the most common bag-of-words-based methods is DBoW2 [galvez2012bags]. In the context of point cloud-based place recognition, BoW3D [cui2022bow3d] has been successfully applied for loop closure on large outdoor trajectories.
@@
For vision-based place recognition, methods such as NetVLAD [arandjelovic2016NetVLADCNNArchitecturea], CosPlace [berton2022RethinkingVisualGeolocalizationa], and MixVPR [ali-bey2023MixVPRFeatureMixinga] use convolutional or transformer-based architectures to predict descriptors. These methods achieve a Top-1 place recognition recall above 0.85 in outdoor urban environments.
For point cloud-based place recognition, PointNetVLAD [uy2018PointNetVLADDeepPoint] extends the NetVLAD architecture to 3D data. Other methods, such as MinkLoc3D [komorowski2021MinkLoc3DPointCloud] and its enhancement [komorowski2022ImprovingPointCloudb], utilize voxelization and sparse convolutions. SVT-Net [fan2022SVTNetSuperLightWeight] further refines this process by incorporating attention mechanisms. These approaches achieve recall rates of 0.87–0.93 in outdoor place recognition for urban environments.
Multimodal place recognition methods have also been developed, integrating image and point cloud data to improve performance. Examples include MinkLoc++ [komorowski2021MinkLocLidarMonoculara] and MSSPlace [mssplace]. In our work, we utilize MinkLoc3D for place recognition as a lightweight and real-time method that relies solely on LiDAR point clouds.
@@ II-C Pose Estimation
For vision-based pose estimation, an example is StereoScan [geiger2011stereoscan], which determines the transformation using stereo image pairs.
For point cloud-based pose estimation, a common approach is Iterative Closest Point (ICP) [besl1992method], which refines the transformation between two point clouds given an initial rough estimate. The RANSAC algorithm [fischler1981random] is often used to generate this initial guess. However, the RANSAC + ICP pipeline is computationally expensive for dense point clouds and performs well only when the input point clouds have significant overlap. According to [muravyev2025prism], this pipeline achieves a recall of 0.75 for point cloud pairs with overlap above 50%, but drops to 0.53 for pairs with overlap above 25%.
Using 2D projections of point clouds (e.g., bird’s-eye views) can improve pose estimation accuracy when overlap is low while also accelerating the matching process compared to full 3D point clouds. Methods such as BVMatch [luo2021bvmatch] and its enhancement BEVPlace++ [luo_BEVPlace_2024] introduce a density-based bird’s-eye-view extraction algorithm and a novel local descriptor, achieving a pose estimation error of 0.4 m.
In our previous work [muravyev2025prism], we introduced a fast bird’s-eye-view matching algorithm based on ORB feature matching with iterative outlier removal. In this work, we enhance our scan matching algorithm for outdoor environments by incorporating curb detection in point clouds and point-to-distance-map optimization.
@@ III PROBLEM STATEMENT
[…]
@@
[…]
$ATE_{mean}=\frac{1}{T^{N}}\sum\limits_{t=1}^{T^{N}}e_{t},$ (2)
$ATE_{median}=Median(e_{1},\dots,e_{T^{N}}).$ (3)
Also we evaluate the localization success rate as the part of time steps where the localization error is lower than a certain threshold:
$SR_{loc}=\frac{1}{T^{N}}\sum\limits_{t=1}^{T^{N}}I(e_{t}<Th).$ (4)
@@ IV METHOD
IV-A Notation
We summarize commonly used notation. Scalars are denoted by lowercase letters (e.g., $t$), vectors by bold lowercase (e.g., $\mathbf{x}$), and matrices by uppercase letters (e.g., $T$). If you already defined notation later, please move the definitive description here.
IV-B Map creation
[…]
@@ IV-C Localization
IV-C1 Overall pipeline
The proposed localization pipeline maintains current state in the graph of locations. At each step $t$, this state is represented as the location $v_{cur}^{t}$ where the robot is currently located, and the robot’s relative position $T_{cur}^{t}$ in this location. The pipeline is similar to the graph maintaining pipeline of the original PRISM-TopoMap method [muravyev2025prism], however, it has no part of a new node addition and checks the localization results for proximity to the current state:
@@ IV-C1 Overall pipeline
  • If we moved outside $v_{cur}^{t-1}$, or our overlapping percent is low, we first try to change $v_{cur}$ to one of its neighbors in the graph (i.e., move along an edge). For this purpose, we use the approximate relative position $((T_{cur}^{t-1})^{-1}(p_{v_{cur}})^{-1}p_{v_{next}})^{-1}$ as an initial guess for scan matching. If we matched successfully to one the neighbors $v_{next}$, then we change $v_{cur}$ to it: $v_{cur}^{t}=v_{next};T_{cur}^{t}=T_{next}$. To prevent false matches, we use a ”jumping threshold” – the maximum distance between an initial guess and the found transform when the matching is considered successful.
@@
  • After all, if we could not change $v_{cur}$ by edges and by the localization results, we change its value to the nearest neighbor in the graph without transform alignment (i.e., try to move along an edge without matching): $v_{cur}^{t}=v_{next};T_{cur}^{t}=((T_{cur}^{t-1})^{-1}(p_{v_{cur}})^{-1}p_{v_{next}})^{-1}$. We do this change if $((T_{cur}^{t-1})^{-1}(p_{v_{cur}})^{-1}p_{v_{next}})^{-1}||<5$ to stay inside the next location. Otherwise, we do not change $v_{cur}$, apply step (1), and consider that the localization is lost.
@@
$\widehat{p_{t}^{next}}=p_{v_{cur}^{t}}^{first}T_{cur}^{t}.$
@@
IV-C2 Place Recognition
For place recognition we use MinkLoc3D [komorowski2021MinkLoc3DPointCloud], selected after training and comparing several models on NCLT — a large $360\deg$ LiDAR dataset close to our domain and to ITLP-Campus [melekhin2024itlp] — where it achieved the best performance (Table I); however, due to noisy GT trajectories, NCLT is unsuitable for strict end-to-end evaluation of the full topological-localization pipeline. For each robot scan, we retrieve the top-5 nearest locations in the topological map by Euclidean distance between MinkLoc3D descriptors and pass these candidates to subsequent pipeline stages.
TABLE I: PR performance on the NCLT [carlevaris2016university] dataset
Method NCLT Inference
R@1 R@5 time, ms
MinkLoc3D [komorowski2021MinkLoc3DPointCloud] 90.5 97.1
ASVT-Net [fan2022SVTNetSuperLightWeight] 86.9 96.0
CSVT-Net [fan2022SVTNetSuperLightWeight] 85.9 95.2
SVT-Net [fan2022SVTNetSuperLightWeight] 81.9 94.6
IV-C3 Scan Matching
@@
To estimate the transform between the reference scan $R$ from the robot, and candidate scan $C$ from the map, we combine three techniques. First, we match the scans using the features $R_{f}$ and $C_{f}$ extracted by ORB detector [rublee2011orb] (for global localization) and by Harris corner feature detector [harris1988combined] for matching along an edge. We find the correspondences between features using FLANN matcher [muja2009flann], as in the PRISM-TopoMap method.
Based on the identified correspondences, we apply a modified RANSAC algorithm to estimate rotation and translation. First, the features from the candidate map are transformed into the reference frame using odometry. For each correspondence, we assign a prior probability $p_{i}$ proportional to $\exp\!\left(-d^{2}/\kappa^{2}\right)$ , where $d$ is the distance between the transformed feature from the candidate map and its corresponding feature in the reference frame and $\kappa=1\ m$. These prior probabilities $p_{i}$ are then used in the sampling step to improve robustness and reduce the number of required iterations.
Figure 4: Scan matching comparison: our previous algorithm (left) [muravyev2025prism] and the proposed algorithm (right). The candidate scan is shown in gray and white, the reference scan is shown in green (wall points) and red (curb points)
An example of the comparison between the proposed scan matching and the PRISM-TopoMap’s one is shown in Fig. 4.